# Algorithms for mean-field variational inference via polyhedral optimization in the Wasserstein space

### v1

@@ 5.3 Approximating the mean-field solution with compatible maps
Let $\rho=\mathcal{N}(0,I)$. For any $\varepsilon>0$, there exists a compatible family $\mathcal{M}$ of optimal transport maps of size $\widetilde{O}(\kappa^{1/2}d^{5/4}/\varepsilon^{1/2})$, with the following property. For any coordinate-wise separable map $\bar{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with Jacobian satisfying the first and second derivative bounds of Theorem 5.4, there exists $\hat{T}\in\underline{\operatorname{cone}}(\mathcal{M};\,\alpha\operatorname{id})_{\sharp}\rho$, with $\alpha=1/\sqrt{L_{V}}$, such that $W_{2}(\bar{T}_{\sharp}\rho,\hat{T}_{\sharp}\rho)=\|\bar{T}-\hat{T}\|_{L^{2}(\rho)}\leq\varepsilon/\ell_{V}^{1/2}$ and $\|D(\bar{T}-\hat{T})\|_{L^{2}(\rho)}\lesssim\kappa^{1/2}d^{1/4}\varepsilon^{1/2}/\ell_{V}^{1/2}$.
@@
The two preceding results show that we can, with prior knowledge of $\pi^{\star}$, construct some $\hat{\pi}_{\diamond}\in\mathcal{P}_{\diamond}$ which is close to $\pi^{\star}$, but it does not guarantee that we can find $\hat{\pi}_{\diamond}$ easily. The next result addresses this issue by showing that $\pi^{\star}$ is close to the minimizer $\pi^{\star}_{\diamond}$ of the KL divergence over $\mathcal{P}_{\diamond}$, and hence can be computed using the algorithms in Section 5.4. As the proof reveals, establishing this statement is related to a geodesic smoothness property for the KL divergence, which is quite non-trivial since the entropy is non-smooth over the full Wasserstein space (see further discussion in the next section). We are able to verify this smoothness property on the geodesic connecting $\hat{\pi}_{\diamond}$ to $\pi^{\star}$ using the bounds on $\|D(\hat{T}_{\diamond}-T^{\star})\|_{L^{2}(\rho)}$ in our approximation results (Theorem 5.6 and Theorem 5.7). The proof of Theorem 5.8 is also found in Section C.3.
@@ 1. $(\Upsilon)$
We remark that when $\mathcal{M}$ is constructed as a direct sum of univariate families via Lemma 3.6 (c.f. Remark 3.10), as in our approximation results (Section 5.3), the matrices $Q$ and $\tilde{Q}$ have a $d\times d$ block diagonal structure; see Section 5.5.1 for details. Consequently, the regularity $\Upsilon$ of the family $\mathcal{M}$ is the same as the regularity parameter for the univariate family used to construct $\mathcal{M}$, and is therefore nominally “dimension-free”. [^4] However, if one wishes to maintain the same quality of approximation in high dimension, our approximation results in Section 5.3 require taking the size of the univariate family to scale mildly with the dimension, and in this case the parameter $\Upsilon$ may indeed scale with the dimension. We leave the detailed quantitative analysis of $\Upsilon$ for future work.
@@
By combining Theorem 5.11 with our approximation result in Theorem 5.8, which provides a bound on $W_{2}(\pi^{\star}_{\diamond},\pi^{\star})$ for explicit choices of $\mathcal{P}_{\diamond}$ with corresponding bounds on the size of $|\mathcal{M}|$, we can then ensure that the iterate $\mu_{(t)}$ is close to $\pi^{\star}$ in the Wasserstein distance. To the best of our knowledge, this constitutes the first accelerated and end-to-end convergence result for mean-field VI. See Section 5.1 for comparisons with the literature.
@@ 5.5.1 Implementation details
For the terms involving explicitly involving $V$, one can draw Monte Carlo samples from the Gaussian $\rho$ and approximate it via empirical averages (assuming access to evaluations of the partial derivatives of $V$).
@@
$DT^{\lambda}(x)=\alpha I_{d}+\operatorname{diag}{\Bigl{(}\sum_{j=1}^{J}\lambda_{j}T_{j}^{\prime}(x_{i})\Bigr{)}_{i=1}^{d}}\,.$
@@
$\int\langle e_{i},(DT^{\lambda})^{-1}(x)\,e_{i}\rangle\,T_{j}^{\prime}(x_{i})\,\rho(\!\,\mathrm{d}x)$ $=\int\frac{T_{j}^{\prime}(x_{i})}{\alpha+\sum_{j^{\prime}=1}^{J}\lambda_{j}T_{j^{\prime}}^{\prime}(x_{i})}\,\rho_{1}(\!\,\mathrm{d}x_{i})\,.$
@@
Again, due to the block diagonal structure of $Q$, this is equivalent to solving $d$ independent projection problems: in each one, we must project a point in $\mathbb{R}^{J}$ onto $\mathbb{R}_{+}^{J}$ in the $Q^{\mathcal{M}_{1}}$-norm. This is a smooth, convex problem that can itself be solved via, e.g., projected gradient descent, or L-BFGS-B (Zhu et al.,, 1997).
@@ 1. $(\Xi)$
    There exists $\Xi>0$ such that for the Gram matrix $Q^{\mathcal{M}_{1}}$ associated with the centered univariate family $\mathcal{M}_{1}$, we have the pointwise bound $\langle Q^{-1},\bar{Q}(x)\rangle\leq\Xi J$ for all $x\in\mathbb{R}$, where $\bar{Q}_{T,\tilde{T}}(x)=\langle T(x),\tilde{T}(x)\rangle$ for $T,\tilde{T}\in\mathcal{M}_{1}$. Here, $J\coloneqq|\mathcal{M}_{1}|$.
@@
Assume that $\pi$ is well-conditioned  $(\mathsf{WC})$ and that $\mathcal{M}$ is regular $(\Upsilon)$ and generated by a univariate family satisfying $(\Xi)$. Let $\pi^{\star}_{\diamond}$ denote the unique minimizer of $\operatorname*{KL}(\cdot\!\;\|\;\!\pi)$ over $\operatorname{cone}(\mathcal{M};\,\alpha\operatorname{id})_{\sharp}\rho$ with $\alpha=1/\sqrt{L_{V}}$. Then, the iterates of stochastic projected gradient descent yield a measure $\mu_{(t)}$ with the guarantee $\sqrt{\ell_{V}}\,W_{2}(\mu_{(t)},\pi^{\star}_{\diamond})\leq\varepsilon$, with a number of iterations bounded by

### v4

@@ 5.3 Approximating the mean-field solution with compatible maps
Let $\rho=\mathcal{N}(0,I)$. For any $\varepsilon>0$, there exists a compatible family $\mathcal{M}$ of optimal transport maps of size $\widetilde{O}(\kappa^{1/2}d^{5/4}/\varepsilon^{1/2})$, with the following property. For any coordinate-wise separable map $\bar{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with Jacobian satisfying the first and second derivative bounds of Theorem 5.4, there exists $\hat{T}\in\underline{\operatorname{cone}}(\mathcal{M};\,\alpha\operatorname{id})$, with $\alpha=1/\sqrt{L_{V}}$, such that $W_{2}(\bar{T}_{\sharp}\rho,\hat{T}_{\sharp}\rho)=\|\bar{T}-\hat{T}\|_{L^{2}(\rho)}\leq\varepsilon/\ell_{V}^{1/2}$ and $\|D(\bar{T}-\hat{T})\|_{L^{2}(\rho)}\lesssim\kappa^{1/2}d^{1/4}\varepsilon^{1/2}/\ell_{V}^{1/2}$.
@@
Remark 5.8.
We note that the worse dependence on $\kappa$ for the smoother dictionary is due to our derivative bounds for the optimal transport map; we have no reason to believe it is fundamental.
The two preceding results show that we can, with prior knowledge of $\pi^{\star}$, construct some $\hat{\pi}_{\diamond}\in\mathcal{P}_{\diamond}$ which is close to $\pi^{\star}$, but it does not guarantee that we can find $\hat{\pi}_{\diamond}$ easily. The next result addresses this issue by showing that $\pi^{\star}$ is close to the minimizer $\pi^{\star}_{\diamond}$ of the KL divergence over $\mathcal{P}_{\diamond}$, and hence can be computed using the algorithms in Section 5.4. As the proof reveals, establishing this statement is related to a geodesic smoothness property for the KL divergence, which is quite non-trivial since the entropy is non-smooth over the full Wasserstein space (see the further discussion in the next section). We are able to verify this smoothness property on the geodesic connecting $\hat{\pi}_{\diamond}$ to $\pi^{\star}$ using the bounds on $\|D(\hat{T}_{\diamond}-T^{\star})\|_{L^{2}(\rho)}$ in our approximation results (Theorem 5.6 and Theorem 5.7). The proof of Theorem 5.9 is also found in Section C.3.
@@ 1. $(\Upsilon)$
We remark that when $\mathcal{M}$ is constructed as a direct sum of univariate families via Lemma 3.6 (c.f. Remark 3.10), as in our approximation results (Section 5.3), the matrices $Q$ and $\tilde{Q}$ have a $d\times d$ block diagonal structure; see Section 5.5.1 for details. Consequently, the regularity $\Upsilon$ of the family $\mathcal{M}$ is the same as the regularity parameter for the univariate family used to construct $\mathcal{M}$, and is therefore nominally “dimension-free”. [^5] However, if one wishes to maintain the same quality of approximation in high dimension, our approximation results in Section 5.3 require taking the size of the univariate family to scale mildly with the dimension, and in this case the parameter $\Upsilon$ may indeed scale with the dimension.
@@
By combining Theorem 5.12 with our approximation result in Theorem 5.9, which provides a bound on $W_{2}(\pi^{\star}_{\diamond},\pi^{\star})$ for explicit choices of $\mathcal{P}_{\diamond}$ with corresponding bounds on the size of $|\mathcal{M}|$, we can then ensure that the iterate $\mu_{(t)}$ is close to $\pi^{\star}$ in the Wasserstein distance. Namely, we ensure that $W_{2}(\mu_{(t)},\pi^{\star})\leq\varepsilon$ provided that we use either of the dictionaries in Theorem 5.9 and we take the number of iterations $t$ as in Theorem 5.12; here, $t$ is the iteration complexity, whereas the bounds on the size of the dictionary in Theorem 5.9 govern the per-iteration cost.
As previously mentioned, our analysis is made possible by bypassing the non-differentiability of entropy over $\mathcal{P}(\mathbb{R})$ and instead optimizing over pointed cones characterized by a univariate compatible family $\mathcal{M}$. Thus, the constant $\Upsilon>0$ should blow up as the polyhedral set approaches $\mathcal{P}(\mathbb{R})$. This fact is summarized in the following lemma for the piecewise-linear family.
Lemma 5.13.
Let $\mathcal{M}$ be the piecewise linear construction of Theorem 5.6. Then, $\Upsilon\lesssim|\mathcal{M}_{1}|^{2}$, where $\mathcal{M}_{1}$ is the generating family in a single dimension. Since $|\mathcal{M}_{1}|=J$, the bound is equivalently $\Upsilon\lesssim J^{2}$.
As a corollary, we can fully characterize the runtime of solving the MFVI problem.
Corollary 5.14 (End-to-end guarantees for MFVI).
Consider the setting of Theorem 5.12. Then the required runtime to compute $\pi_{\diamond}^{\star}$ becomes
$t=O(J\kappa^{1/2}\log(\sqrt{\kappa}d/\varepsilon))\,.$
If $\pi_{\diamond}^{\star}$ is meant to approximation $\pi^{\star}$, then we can use the approximation guarantees from Theorem 5.9 to obtain the complete convergence guarantee of
$t=O({\kappa^{5/2}}d^{1/2}\varepsilon^{-1}\log(\sqrt{\kappa}d/\varepsilon))\,.$
As we demonstrate below, the regime of $J=O(1)$ appears to suffice numerically. To the best of our knowledge, this constitutes the first accelerated and end-to-end convergence result for mean-field VI. See Section 5.1 for comparisons with the literature.
@@ 5.5.1 Implementation details
For the terms explicitly involving $V$, one can draw Monte Carlo samples from the Gaussian $\rho$ and approximate them via empirical averages (assuming access to evaluations of the partial derivatives of $V$).
@@
$DT^{\lambda}(x)=\alpha I_{d}+\operatorname{diag}{\Bigl{(}\sum_{j=1}^{J}\lambda_{i,j}T_{j}^{\prime}(x_{i})\Bigr{)}_{i=1}^{d}}\,.$
@@
$\int\langle e_{i},(DT^{\lambda})^{-1}(x)\,e_{i}\rangle\,T_{j}^{\prime}(x_{i})\,\rho(\!\,\mathrm{d}x)$ $=\int\frac{T_{j}^{\prime}(x_{i})}{\alpha+\sum_{j^{\prime}=1}^{J}\lambda_{i,j^{\prime}}T_{j^{\prime}}^{\prime}(x_{i})}\,\rho_{1}(\!\,\mathrm{d}x_{i})\,.$
@@
Again, due to the block diagonal structure of $Q$, this is equivalent to solving $d$ independent projection problems: in each one, we must project a point in $\mathbb{R}^{J}$ onto $\mathbb{R}_{+}^{J}$ in the $Q^{\mathcal{M}_{1}}$-norm. This is a smooth, convex problem that can itself be solved via, e.g., projected gradient descent, or L-BFGS-B (Zhu et al.,, 1997), or any standard quadratic program solver.
@@ 1. $(\Xi)$
    There exists $\Xi>0$ such that for the Gram matrix $Q^{\mathcal{M}_{1}}$ associated with the centered univariate family $\mathcal{M}_{1}$, we have the pointwise bound $\langle Q^{-1},\bar{Q}(x)\rangle\leq\Xi J$ for all $x\in\mathbb{R}$, where $\bar{Q}_{T,\tilde{T}}(x)={T(x)\tilde{T}(x)}$ for $T,\tilde{T}\in\mathcal{M}_{1}$. Here, $J\coloneqq|\mathcal{M}_{1}|$.
Similarly to Lemma 5.13, we can also quantify $\Xi$ for the piecewise linear dictionary.
Lemma 5.15.
Let $\mathcal{M}$ be the piecewise linear construction of Theorem 5.6. Then, $\Xi\lesssim|\mathcal{M}_{1}|^{2}$, where $\mathcal{M}_{1}$ is the generating family in a single dimension. Since $|\mathcal{M}_{1}|=J$, the bound is equivalently $\Xi\lesssim J^{2}$.
@@
Assume that $\pi$ is well-conditioned  $(\mathsf{WC})$ and that $\mathcal{M}$ is regular $(\Upsilon)$ and generated by a univariate family satisfying $(\Xi)$. Let $\pi^{\star}_{\diamond}$ denote the unique minimizer of $\operatorname*{KL}(\cdot\!\;\|\;\!\pi)$ over $\operatorname{cone}(\mathcal{M};\,\alpha\operatorname{id})_{\sharp}\rho$ with $\alpha=1/\sqrt{L_{V}}$. Then, for all sufficiently small $\varepsilon$, the iterates of stochastic projected gradient descent yield a measure $\mu_{(t)}$ with the guarantee $\sqrt{\ell_{V}}\,\mathbb{E}[W_{2}(\mu_{(t)},\pi^{\star}_{\diamond})]\leq\varepsilon$, with a number of iterations bounded by
@@
As with Theorem 5.12, we can state the following corollary given that $\Xi$ is uniformly bounded via Lemma 5.15.
Corollary 5.18 (End-to-end guarantees with SPGD).
Consider the setting of Theorem 5.17. Then the required runtime to estimate $\pi_{\diamond}^{\star}$ becomes
$t=O(dJ^{3}\kappa^{2}\varepsilon^{-2}\log(\sqrt{\ell_{V}}d/\varepsilon))\,.$
If $\pi_{\diamond}^{\star}$ is meant to approximation $\pi^{\star}$, then we can use the approximation guarantees from Theorem 5.9 to obtain the complete convergence guarantee of
$t=O({\kappa^{8}}d^{5/2}\varepsilon^{-5}\log(\sqrt{\ell_{V}}d/\varepsilon))\,.$
6 Numerical experiments
Figure 1: KDEs for the optimal product Gaussian mixture and our algorithm.
Figure 2: Our algorithm is robust to the choice of $\alpha$.
We showcase our proposed MFVI algorithm on numerical experiments. Experimental details are deferred to Appendix D, and the code to reproduce the experiments is available here. Across all experiments, which include low- and high-dimensional settings, we took the piecewise linear dictionary (Theorem 5.6) with the same value for the size $J=|\mathcal{M}_{1}|=28$ of the univariate family (hence $|\mathcal{M}|=Jd$), and we ran stochastic gradient descent (without acceleration) with a batch size of 2000 samples per iteration.
6.1 Product Gaussian mixture
In our first experiment, the target is a mixture of four Gaussians in $\mathbb{R}^{2}$ which is itself a product measure. Despite the non-log-concavity, our algorithm correctly recovers the correct target. Though, we note that this approach is sensitive to the initialization, but this is expected as the landscape is non-convex.
6.2 Non-isotropic Gaussian
Next, we computed the mean-field approximation of a randomly generated centered and non-isotropic Gaussian in dimension $d=5$. Letting $\Sigma$ denote the covariance matrix, the mean-field approximation is also a Gaussian with diagonal covariance and entries $(\Sigma_{\text{MF}})_{i,i}=1/(\Sigma^{-1})_{i,i}$ (see Section D.2 for a calculation of this fact).
In Figure 2, we plot the $W_{2}^{2}$ error between the covariance matrix of our algorithm iterate (computed from samples) and $\Sigma_{\rm MF}$, which is a lower bound on the true $W_{2}^{2}$ distance (cf. Cuesta-Albertos et al.,, 1996).
In this case, the optimal parameter choice $\alpha^{*}$ is known, though this is rarely the case in practice. We ran our algorithm for various choices of $\alpha$, fixing all other parameters to be the same. We see that our algorithm does not depend heavily on the choice of hyperparameter $\alpha$, and the practitioner can safely choose a small value of $\alpha$ without sacrificing performance.
6.3 Synthetic Bayesian logistic regression
As a final example, we turn to Bayesian logistic regression on synthetic data; precise details are deferred to Section D.3. In summary, we are given i.i.d. data $(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$ for $i=1,\ldots,n$, (where $d=20$ and $n=100$) from which we want to recover a parameter $\theta$. When assuming an improper (Lebesgue) prior on $\theta$, the posterior is given by
$V(\theta)=\sum_{i=1}^{n}\bigl{[}\log(1+\exp(\theta^{\top}X_{i}))-Y_{i}\,\theta^{\top}X_{i}\bigr{]}\,.$
Note that $V$ is not strongly convex as $V(\theta)$ behaves like a linear function as $\|\theta\|\to+\infty$. With $V$ and $\nabla V$ in hand, our algorithm is fully implementable. As we considered an improper prior, a comparison to CAVI is not possible. Instead, we compared against standard Langevin Monte Carlo (LMC). The final histograms were generated using 2000 samples from both the mean-field VI algorithm and LMC. Figure 3 contains the 20 marginals for both our approach and LMC, which are closely aligned.
Figure 3: Histograms of the first ten marginals computed via our mean-field VI algorithm vs. Langevin Monte Carlo for a $20$-dimensional Bayesian logistic regression example.
7 Extension to mixtures of product measures
In this section, we extend our methodology to approximations via mixtures of product measures. The motivation is simply that many more measures can be approximated via mixtures of (approximate) product measures, e.g., Gibbs distributions with small gradient complexity (Eldan,, 2018; Eldan and Gross,, 2018; Austin,, 2019; Jain et al.,, 2019).
In Section 5.4, we minimized $\operatorname*{KL}(\cdot\!\;\|\;\!\pi)$ over $\underline{\operatorname{cone}}(\mathcal{M};\,\alpha\operatorname{id})_{\sharp}\rho$, where $\underline{\operatorname{cone}}(\mathcal{M};\,\alpha\operatorname{id})$ is parameterized by the pair $(\lambda,v)\in\mathcal{M}\coloneqq\mathbb{R}_{+}^{|\mathcal{M}|}\times\mathbb{R}^{d}$, equipped with the norm $\left\|\cdot\right\|_{Q\oplus I_{d}}$. In this section, following Lambert et al., (2022), a mixture of product measures is specified by a mixing measure $P\in\mathcal{P}(\mathcal{M})$ and corresponds to the measure $\mu_{P}\coloneqq\int(T^{\lambda,v})_{\sharp}\rho\,P(\!\,\mathrm{d}\lambda,\!\,\mathrm{d}v)$. We can now equip the space $\mathcal{P}(\mathcal{M})$ with the Wasserstein geometry (with respect to $\left\|\cdot\right\|_{Q\oplus I_{d}}$), and we shall derive the Wasserstein gradient flow of the functional $P\mapsto\operatorname*{KL}(\mu_{P}\!\;\|\;\!\pi)$.
This approach to mixture modelling is inspired by the distance on Gaussian mixtures proposed in Chen et al., (2019); Delon and Desolneux, (2020); see Bing et al., (2023) for a statistical perspective.
In this section, we again use the abstract parameterization $T^{\lambda,v}=\alpha\operatorname{id}+\sum_{T\in\mathcal{M}}\lambda_{T}T+v$. Proofs are given in Appendix E.
Theorem 7.1.
The Wasserstein gradient flow of $P\mapsto\operatorname*{KL}(\mu_{P}\!\;\|\;\!\pi)$ is the flow ${(P^{(t)})}_{t\geq 0}$ specified as follows. For each $t\geq 0$, $P^{(t)}$ is the law of $(\lambda^{(t)},v^{(t)})$, where
$\dot{\lambda}^{(t)}_{T}$ $=-\int\bigl{\langle}\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)},v^{(t)}},T\bigr{\rangle}\,\mathrm{d}\rho\,,$ $\text{for}\leavevmode\nobreak\ T\in\mathcal{M}\,,$
$\dot{v}^{(t)}$ $=-\int\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)},v^{(t)}}\,\mathrm{d}\rho\,.$
In practice, we use a finite number $K$ of mixture components, in which case
$P=\frac{1}{K}\sum_{k=1}^{K}\delta_{(\lambda[k],v[k])}\,,\qquad\mu_{P}=\frac{1}{K}\sum_{k=1}^{K}(T^{\lambda[k],v[k]})_{\sharp}\rho\,.$ (19)
The system of ODEs above then becomes an interacting particle system:
$\dot{\lambda}_{T}^{(t)}[k]$ $=-\int\bigl{\langle}\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)}[k],v^{(t)}[k]},T\bigr{\rangle}\,\mathrm{d}\rho\,,$ $\text{for}\leavevmode\nobreak\ T\in\mathcal{M}\,,$
$\dot{v}^{(t)}[k]$ $=-\int\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)}[k],v^{(t)}[k]}\,\mathrm{d}\rho\,.$
The particles interact through the common term $\log\mu_{P^{(t)}}$. More explicitly, by the change of variables formula,
$\mu_{P}=\frac{1}{K}\sum_{k=1}^{K}\frac{\rho\circ(T^{\lambda[k],v[k]})^{-1}}{\det DT^{\lambda[k],v[k]}\circ(T^{\lambda[k],v[k]})^{-1}}\,.$
Note that computing $\nabla\log\mu_{P}$ now requires taking the second derivative of the transport maps, which hinders implementation. In this case, a smooth family $\mathcal{M}$ is required.
The dynamics (19) maintains equal weights for each of the particles at each time. We can similarly derive the gradient flow with respect to the Wasserstein–Fisher–Rao (or Hellinger–Kantorovich) geometry, which captures unbalanced optimal transport (Liero et al.,, 2016; Chizat et al.,, 2018; Liero et al.,, 2018). The use of this geometry for sampling was pioneered in Lu et al., (2019).
Theorem 7.2.
The Wasserstein–Fisher–Rao gradient flow of $P\mapsto\operatorname*{KL}(\mu_{P}\!\;\|\;\!\pi)$, initialized at $P^{(0)}=\sum_{k=1}^{K}w^{(0)}[k]\,\delta_{(\lambda^{(0)}[k],v^{(0)}[k])}$ with $\sum_{k=1}^{K}w^{(0)}[k]=1$, can be described as follows. For each time $t\geq 0$, let $P^{(t)}=\sum_{k=1}^{K}w^{(t)}[k]\,\delta_{(\lambda^{(t)}[k],v^{(t)}[k])}$ and $r^{(t)}[k]\coloneqq\sqrt{w^{(t)}[k]}$. Then,
$\dot{\lambda}_{T}^{(t)}[k]$ $=-\int\bigl{\langle}\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)}[k],v^{(t)}[k]},T\bigr{\rangle}\,\mathrm{d}\rho\,,$ $\text{for}\leavevmode\nobreak\ T\in\mathcal{M}\,,$
$\dot{v}^{(t)}[k]$ $=-\int\nabla\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)}[k],v^{(t)}[k]}\,\mathrm{d}\rho\,,$
$\dot{r}^{(t)}[k]$ $=-\Bigl{(}\int\log\frac{\mu_{P^{(t)}}}{\pi}\circ T^{\lambda^{(t)}[k],v^{(t)}[k]}\,\mathrm{d}\rho-\int\log\frac{\mu_{P^{(t)}}}{\pi}\,\mathrm{d}\mu_{P^{(t)}}\Bigr{)}\,r^{(t)}[k]\,.$
We leave it as an open question to obtain convergence rates for this flow.